There are four shells A, B, C and D of radii a, 2a, 3a and 4a, respectively. Shells B and D are given charges +q and -q, respectively. Shell C is now earthed. The potential difference between A and C is
  • Kq/2a
  • Kq/3a
  • Kq/4a
  • Kq/6a

Grade:11

1 Answers

abdul
4 years ago
VA=k(0/a + Q/2a + 0/3a – Q/4a)
VA=kQ/4a
 
VB=K(0/a +Q/3a +0/3a -Q/4a)
VB=KQ/12a
 
va-vb=kQ/4a – kQ/12a= k q /6a